OpenStudy (anonymous):

Bonnie and Clyde are sliding a 300 kg bank safe across the floor to their getaway car. The safe slides with a constant speed if Clyde pushes from behind with 385 N of force while Bonnie pulls forward on a rope with 350 N of force. What is the safe’s coefficient of kinetic friction on the bank floor?

OpenStudy (xishem):

Start by drawing a force diagram.

OpenStudy (anonymous):

I already did, and I come up with an answer of 0.148 (u = 0.148) but it is wrong, I don't know why.

OpenStudy (xishem):

\[\sum F_y=F_N-F_g=0\]\[\sum F_x=385N+350N-f_{fric}=0\]Is this consistent with your equations? I assumed the safe was being pushed/pulled to the right.

OpenStudy (anonymous):